OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x + 3y = –12

OpenStudy (texaschic101):

to find the y intercept, sub in 0 for x and solve for y 2x + 3y = -12 2(0) + 3y = -12 3y = -12 y = -4.....so your y intercept is (0,-4) to find the x intercept, sub in 0 for y and solve for x 2x + 3y = -12 2x + 3(0) = -12 2x = -12 x = -6...the x intercept is (-6,0) any questions ?

OpenStudy (anonymous):

No! thank you so much :)

OpenStudy (anonymous):

The x-intercept is where y is equal to 0. So to find the x-intercept, plug in 0 for y, and you get 2x =- 12. The only way that works is if x is 6, so that's your intercept. The point would be (-6, 0).

OpenStudy (texaschic101):

anytime :)